# MERLIN: Multi-stagE query performance prediction for dynamic paRallel oLap pIpeliNe

### v1

@@ 6.2. QPP Accuracy
MERLIN 1.46 1.49 1.90 1.98 2.05 2.49 1.84 1.73 2.78 2.88 3.22 3.84
MERLIN w/o OCP 1.70 1.61 2.46 3.13 3.26 3.35 2.05 1.57 3.62 5.26 8.27 9.49
MERLIN w/o ResAttn 2.09 1.97 3.05 3.25 4.29 6.52 2.48 1.90 4.95 5.62 6.68 7.72
Figure 5. Q-Error distribution comparison.
@@
The results of MERLIN’s and baselines’ accuracy are shown in Table 2 and Figure 5. On both TPC-H and TPC-DS workloads, MERLIN significantly outperforms all baselines. The Q-Error distribution shows consistent performance gains across OLAP workloads of varying size and complexity. The TCN’s Q-Error is quite low on TPC-H but increases a lot on TPC-DS, especially the max Q-Error. We also noticed that both GCN and TCN can reach similar accuracy compared to MERLIN on TPC-H, but their Q-Error gains a lot on TPC-DS. This proves the effectiveness of these two widely used methods. But it also indicates that when it comes to complex OLAP workloads on high-performance DBMSs, such methods lose their accuracy due to a lack of targeted design. On the other hand, MERLIN maintains a stable lead on both workloads. The QPPNet’s long-tail distribution problem on TPC-H is indeed serious. We have tried everything that we could such as adjusting its hyper-parameters and switching to better loss functions like LeakyReLU and Swish. This is the best result we can get. We also noticed that, although TPC-DS is a much more complex workload, the max Q-Error of QPPNet was significantly reduced compared to TPC-H.
@@
We think the problem is that QPPNet is not a resource-aware approach that cannot model parallel pipeline execution well in certain situations. With enough training templates, it may perform better due to the ability to generalize. However, TPC-H is a workload that only has a few templates, 23.5% of templates used for testing is too high for them. Therefore, when there are serious resource competitions in some queries, QPPNet which is trained on limited TPC-H workloads cannot predict the latency accurately. This result provides indirect evidence that traditional learning methods are difficult to adapt to high-performance OLAP DBMSs and the resource-aware mechanism is crucial in OLAP DBMS like ClickHouse.
@@
In order to more intuitively compare the error distribution between different methods, we grouped the methods by latency range and plotted the accuracy box plots for each method since the latency can reflect the complexity of the query. The results are shown in Figure 6. The results show that the Q-Error of the QPP task in different query execution time groups does not show an obvious increasing trend. We think this is because the accuracy of query-driven QPP methods is more affected by the generalization between different templates than by the size of the query plan. The long tail distribution of methods such as QPPNet and TCN is concentrated in a certain group interval because the execution time of queries generated by the same query template is more concentrated than that of queries generated by different templates. This phenomenon occurs when the QPP method has poor generalization performance on individual templates. Taking the QPPNet as an example, its Q-Error has a significant peak at the 25-50th group. As analyzed above, the number of training templates is too few and the testing templates’ ratio is too high, so QPPNet cannot handle resource competition between operators well on individual templates. So most of the queries generated by the specific template fall into the same interval and cause this result. For resource-aware methods like MERLIN and RAAL, a relatively stable Q-Error can be maintained under different query complexities, that is, their long-tail distribution problem is relatively mild. This reveals that in high-performance OLAP DBMS such as ClickHouse, the usage of various hardware resources in the system and the resource competition of operators are crucial to the accuracy of QPP tasks.
@@
Ablation Study. By comparing the accuracy of ablation methods MERLIN w/o OCP and MERLIN w/o ResAttn, we can conclude that the resource attention mechanism has the biggest improvement in the prediction accuracy and both modules can significantly improve the accuracy.
@@
Figure 6. Q-Error distribution grouped by queries’ latency. Box plots are drawn in the order of the legend.
6.3. Performance Comparison
@@
The training speed, inference latency, and model size are crucial to the practicality of the QPP approach. We report the performance of MERLIN and compare it with baselines as shown in Table 3. Since TCN is a component of MERLIN we didn’t include it for comparison. For the performance of training speed, the QueryFormer is the fastest because it uses a well-optimized Transformer architecture. Since MERLIN employs multi-stage models and dynamic data-flow tree expression, which involves a re-index of tensors. So MERLIN is a little bit slower than QueryFormer and becomes the second fastest method among all compared ones. QPPNet suffers from its dynamic network structure, significantly slowing down its training speed. RAAL uses the most complex network structure and training takes 599.5 seconds per epoch. The training speed itself can not reflect the actual training efficiency, so we draw their convergence curve in Figure 7. The QPPNet convergences the fastest, which takes 5 epochs to drop from 2.14 to 1.89. But its average Q-Error is quite high compared with other methods. Its accuracy on the test set shows similar results. However, as the methods that have the lowest Q-Error on the train set, RAAL and MERLIN w/o OCP have higher Q-Error than MERLIN on the test set, which shows their poorer generalization ability. In general, most methods reach their lowest Q-Error at around the 14th epoch. The difference in convergence is much insignificant than their training time cost per epoch.
@@
In terms of model size and inference speed, we summarized each approach’s model size by counting each parameter’s size obtained by the state_dict(). Note that all OCP models’ parameters are counted in. As shown in Table 3, MERLIN has the smallest model and fastest inference latency among all methods. RAAL suffers from its recurrent neural network mechanism that can not perform parallel inference and has the highest inference latency. QPPNet needs to build very large networks to match the complex OLAP queries, thus it has the highest inference latency.
@@
Table 3. Performance comparison of baselines, MERLIN. The data is models trained on TPC-H workload.
@@
MERLIN 0.094 78.4 4.2
Figure 7. Q-Error with a standard error during training.
6.4. Robustness to Cardinality Errors
The cardinality errors usually have a significant impact on cost estimation and query performance prediction tasks, as the input scale directly decides the cost. Therefore, we evaluate MERLIN’s robustness to cardinality errors in this subsection. Due to the complexity of ClickHouse’s pipeline builder and dynamic executor, it is hard to adapt multiple cardinality estimation methods to its kernel. Instead, we generate Q-Errors that follow a log-normal distribution and simulate the estimation result by multiplying it with the true cardinality collected by MERLIN’s Runtime Tracker. As shown in Figure 8, the Q-Error distribution of generated cardinality is similar to real estimation with a long-tailed distribution.
Figure 8. Distribution of generated Q-Error with log-scale x-axis. The Q-Error’s maximum value is 740.08.
We assess MERLIN’s accuracy changing as the changing of the Q-Error distribution on TPC-H workloads. Due to QPPNet’s accuracy being significantly lower than other baselines, we did not include it in the experiment. We set log-normal distribution’s $\mu=0$ and the standard deviation $\sigma=\{1.0,1.5\}$.
Table 4. Robustness of Cardinality Error, measured by mean Q-Error.
@@ 6.4. Robustness to Cardinality Errors
MERLIN 1.46 1.49 2.59
Figure 9. Robustness Q-Error distribution comparison.
As Table 4 and Figure 9 show, the predicted query performance’s mean and max Q-Error increase as the $\sigma$ of Q-Error distribution increases as we expected. In terms of max Q-Error, MERLIN outperforms baselines on each setting of $\sigma$. Although QueryFormer’s mean Q-Error almost doesn’t increase, its max Q-Error increases much faster than MERLIN and RAAL. But in terms of mean Q-Error, MERLIN’s is higher than the two baselines when $\sigma=1.5$. This shows that although MERLIN continues to benefit from the insignificant long-tail distribution caused by its design, its robustness remains to be improved under larger cardinality estimation errors.
7. Conclusion and Future Works
In this paper, we proposed a novel query performance prediction approach MERLIN for high-performance OLAP DBMSs with SIMD operators and dynamic parallel pipeline. By carefully tracking and collecting data with the Runtime Tracker and employing a multi-stage resource-competition-aware model, MERLIN can capture the resource competition and dynamic pipeline modifications better in such DBMSs. We implemented MERLIN for ClickHouse and adapted three baselines for comparison, the experimental results show that MERLIN outperforms all baselines in accuracy, model size, and inference speed, while its training speed and robustness are also practical. We plan to expand MERLIN and create a cross-DBMS query performance prediction plugin that can automatically adapt different DBMSs.

### v2

@@ VI-B QPP Accuracy
CONCERTO 1.46 1.49 1.90 1.98 2.05 2.49 1.84 1.73 2.78 2.88 3.22 3.84
CONCERTO w/o OCP 1.70 1.61 2.46 3.13 3.26 3.35 2.05 1.57 3.62 5.26 8.27 9.49
CONCERTO w/o ResAttn 2.09 1.97 3.05 3.25 4.29 6.52 2.48 1.90 4.95 5.62 6.68 7.72
Figure 5: Q-Error distribution comparison.
@@
The results of CONCERTO’s and baselines’ accuracy are shown in Table II and Figure 5. On both TPC-H and TPC-DS workloads, CONCERTO significantly outperforms all baselines. The Q-Error distribution shows consistent performance gains across OLAP workloads of varying size and complexity. The TCN’s Q-Error is quite low on TPC-H but increases a lot on TPC-DS, especially the max Q-Error. We also noticed that both GCN and TCN can reach similar accuracy compared to CONCERTO on TPC-H, but their Q-Error gains a lot on TPC-DS. This proves the effectiveness of these two widely used methods. But it also indicates that when it comes to complex OLAP workloads on high-performance DBMSs, such methods lose their accuracy due to a lack of targeted design. On the other hand, CONCERTO maintains a stable lead on both workloads. The QPPNet’s long-tail distribution problem on TPC-H is indeed serious. We have tried everything that we could, such as adjusting its hyperparameters and switching to better loss functions like LeakyReLU and Swish. This is the best result we can get. We also noticed that, although TPC-DS is a much more complex workload, the max Q-Error of QPPNet was significantly reduced compared to TPC-H.
@@
We think the problem is that QPPNet is not a resource-aware approach that cannot model parallel pipeline execution well in certain situations. With enough training templates, it may perform better due to the ability to generalize. However, TPC-H is a workload that only has a few templates, and 23.5% templates used for testing is too high for them. Therefore, when there are serious resource competitions in some queries, QPPNet, which is trained on limited TPC-H workloads, cannot predict the latency accurately. This result provides indirect evidence that traditional learning methods are difficult to adapt to high-performance OLAP DBMSs, and the resource-aware mechanism is crucial in them.
@@
In order to more intuitively compare the error distribution between different methods, we grouped the methods by latency range and plotted the accuracy box plots for each method, since the latency can reflect the complexity of the query. The results are shown in Figure 6. The results show that the Q-Error of the QPP task in different query execution time groups does not show an obvious increasing trend. We think this is because the accuracy of query-driven QPP methods is more affected by the generalization between different templates than by the size of the query plan. The long tail distribution of methods such as QPPNet and TCN is concentrated in a certain group interval because the execution time of queries generated by the same query template is more concentrated than that of queries generated by different templates. This phenomenon occurs when the QPP method has poor generalization performance on individual templates. Taking the QPPNet as an example, its Q-Error has a significant peak at the 25- 50th group. As analyzed above, the number of training templates is too few and the testing templates’ ratio is too high, so QPPNet cannot handle resource competition between operators well on individual templates. So, most of the queries generated by the specific template fall into the same interval and cause this result. For resource-aware methods like CONCERTO and RAAL, a relatively stable Q-Error can be maintained under different query complexities, that is, their long-tail distribution problem is relatively mild. This reveals that in high-performance OLAP DBMS, the usage of various hardware resources in the system and the resource competition of operators are crucial to the accuracy of QPP tasks.
@@
Ablation Study. By comparing the accuracy of ablation methods CONCERTO w/o OCP and CONCERTO w/o ResAttn, we can conclude that the resource attention mechanism has the biggest improvement in the prediction accuracy, and both modules can significantly improve the accuracy.
@@
Figure 6: Q-Error distribution grouped by queries’ latency. Box plots are drawn in the order of the legend.
VI-C Performance Comparison
@@
The training speed, inference latency, and model size are crucial to the practicality of the QPP approach. We report the performance of CONCERTO and compare it with baselines as shown in Table III. Since TCN is a component of CONCERTO, we didn’t include it for comparison. For the performance of training speed, the QueryFormer is the fastest because it uses a well-optimized Transformer architecture. Since CONCERTO employs multi-stage models and dynamic data-flow tree expression, which involves a re-indexing of tensors. So CONCERTO is a little bit slower than QueryFormer and becomes the second fastest method among all compared ones. QPPNet suffers from its dynamic network structure, significantly slowing down its training speed. RAAL uses the most complex network structure, and training takes 599.5 seconds per epoch. The training speed itself can not reflect the actual training efficiency, so we draw their convergence curve in Figure 7. The QPPNet converges the fastest, which takes 5 epochs to drop from 2.14 to 1.89. But its average Q-Error is quite high compared with other methods. Its accuracy on the test set shows similar results. However, as the methods that have the lowest Q-Error on the train set, RAAL and CONCERTO w/o OCP have higher Q-Error than CONCERTO on the test set, which shows their poorer generalization ability. In general, most methods reach their lowest Q-Error at around the 14th epoch. The difference in convergence is much significant than their training time cost per epoch.
@@
In terms of model size and inference speed, we summarized each approach’s model size by counting each parameter’s size obtained by the state_dict(). Note that all OCP models’ parameters are counted in. As shown in Table III, CONCERTO has the smallest model and fastest inference latency among all methods. RAAL suffers from its recurrent neural network mechanism, which can not perform parallel inference and has the highest inference latency. QPPNet needs to build very large networks to match the complex OLAP queries, thus, it has the highest inference latency.
@@
TABLE III: Performance comparison of baselines, CONCERTO. The data is from models trained on the TPC-H workload.
@@
CONCERTO 0.094 78.4 4.2
Figure 7: Q-Error with a standard error during training.
VI-D Robustness to Cardinality Errors
The cardinality errors usually have a significant impact on cost estimation and query performance prediction tasks, as the input scale directly decides the cost. Therefore, we evaluate CONCERTO’s robustness to cardinality errors in this subsection. Due to the complexity of ClickHouse’s pipeline builder and dynamic executor, it is hard to adapt multiple cardinality estimation methods to its kernel. Instead, we generate Q-Errors that follow a log-normal distribution and simulate the estimation result by multiplying it with the true cardinality collected by CONCERTO’s Runtime Tracker. As shown in Figure 8, the Q-Error distribution of generated cardinality is similar to real estimation with a long-tailed distribution.
Figure 8: Distribution of generated Q-Error with log-scale x-axis. The Q-Error’s maximum value is 740.08.
We assess CONCERTO’s accuracy as the change of the Q-Error distribution on TPC-H workloads. Due to QPPNet’s accuracy being significantly lower than other baselines, we did not include it in the experiment. We set log-normal distribution’s $\mu=0$ and the standard deviation $\sigma=\{1.0,1.5\}$.
TABLE IV: Robustness of Cardinality Error, measured by mean Q-Error.
@@ VI-D Robustness to Cardinality Errors
CONCERTO 1.46 1.49 2.59
Figure 9: Robustness Q-Error distribution comparison.
As Table IV and Figure 9 show, the predicted query performance’s mean and max Q-Error increase as the $\sigma$ of Q-Error distribution increases, as we expected. In terms of max Q-Error, CONCERTO outperforms baselines on each setting of $\sigma$. Although QueryFormer’s mean Q-Error almost doesn’t increase, its max Q-Error increases much faster than CONCERTO and RAAL. But in terms of mean Q-Error, CONCERTO’s is higher than the two baselines when $\sigma=1.5$. This shows that although CONCERTO continues to benefit from the insignificant long-tail distribution caused by its design, its robustness remains to be improved under larger cardinality estimation errors.
VII Conclusion and Future Works
In this paper, we propose a novel query performance prediction approach, CONCERTO, for high-performance OLAP DBMSs with complex query execution mechanisms like vectorized operators, DAG execution plan, and dynamic parallel pipeline. By carefully tracking and collecting data with the Runtime Tracker and employing a multi-stage resource-competition-aware model, CONCERTO can capture the resource competition and dynamic pipeline modifications better in such DBMSs. We implemented CONCERTO for ClickHouse and adapted three baselines for comparison, The experimental results show that CONCERTO outperforms all baselines in accuracy, model size, and inference speed, while its training speed and robustness are also practical. We plan to expand CONCERTO and create a cross-DBMS query performance prediction plugin that can automatically adapt to different DBMSs.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (NSFC) (62232005).